OpenStudy (anonymous):

a boat is 144 feet from the base of cliff. If the distance from the top of the cliff to the boat is 36 more than twice the height of the cliff. find the height of the cliff. find the height of the cliff. round to the nearest tenth of a foot if necessary.

OpenStudy (bahrom7893):

z = 36+2y

OpenStudy (bahrom7893):

x = 144

OpenStudy (bahrom7893):

x^2 + y^2 = z^2

OpenStudy (bahrom7893):

3 equations and 3 unknowns

OpenStudy (campbell_st):

draw a picture  now use pythagoras's theorem

OpenStudy (campbell_st):

so you will have \[x^2 + 144^2 = (2x + 36)^2\] its just a case of solving for x